OpenStudy (jrmeyer):

When x is added to the numerator and subtracted from the denominator of the fraction a/b, where a <> b and b <> 0, the value of the fraction is changed to c/d. Find x

OpenStudy (anonymous):

(a+x)/(b-x) = c/d d(a+x)=c(b-x) ad+dx=bc-cx (c+d)x=ad+bc x=(ad+bc)/(c+d)

OpenStudy (anonymous):

Did u mean that?

OpenStudy (jrmeyer):

Thank you!